Mrs. Robinson is 75-years-old and weighs 110 kg. You attempt to determine the optimal maintenance dose. Your target serum concentration is 54.7 ug/ml, bioavailability is 0.86, CL = 19 ml/H/kg, and Vd = 1.6 L/kg. The standard dosing interval for this drug is 12 hours.
Answer #1

Peak concentration = 54.7 microgram / ml

= 0.0547 mg/ ml = 0.0547× 0.001 = 0.0000547mg/L

Clearance of drug in the body = 19 ml /h/kg

= 0.019L/ h/ kg

Weight of the patient is 110 kg

Clearance = 0.019×110= 2.09L/ h

Bioavailability = 0.86

Maintenance dose = peak concentration × clearance / bioavailability

= 0.0000547(mg/L) × 2.09(L/h) ÷0.86

= 0.000132 mg/ h